A cyclist traveled
10
kilometers per hour faster than an in-line skater. In the time it took the cyclist to travel
40
kilometers, the skater had gone
20
kilometers. Find the speed of the skater.

Answer by jorel555(1290)

Let the speed of the skater be n. Then:
20/n=40/n+10
20(n+10)=40n
20n=200
n=10
The skaters speed was 10 kmh!!!!!!!!